Groupoids and Stacks in Physics and Geometry

June 28 – July 2, 2004 CIRM Luminy Schedule

Monday, June 28, Morning Session

9:00 Yvette Kosmann- The infinitesimals of Lie groupoids:
–9:50 Schwarzbach  Lie algebroids
10:00 Lawrence Breen Gerbes and 2-groupoids

–10:50

11:00 Coffee
11:30 Kirill MacKenzie Gerbes avant la lettre

–12:20 Afternoon Session

4:30 Coffee
5:00 Peter Bouwknegt T-duality and topology change

–5:50

6:00 Eric Sharpe  Some progress in the physics of
–6:50  string compactifications on stacks

Tuesday, June 29, Morning Session

9:00 Jean Renault  Continuous trace algebras and the
–9:50  Dixmier-Douady invariant
10:00 Ieke Moerdijk  Gerbes, Torsors, and the classification of
–10:50  regular Lie groupoids
11:00 Coffee
11:30 John Lott  Higher-degree alalogs of the determinant line bundle

–12:20 Afternoon Session

4:30 Coffee
5:00 Chenchang Zhu  Integrating Lie algebroids via stacks

–5:50

6:00 Alan Weinstein  Integration of complex Lie algebroids

–6:50 Wednesday, June 30, Morning Session

9:00 Eckhard Meinrenken Equivariant cohomology and
–9:50  the Maurer-Cartan equation
10:00 Clemens Berger  Higher groupoids and homotopy type

–10:50

11:00 Coffee
11:30 Ezra Getzler  L_{∞}-algebras and n-groupoids

–12:20 No Afternoon Session

Thursday, July 1, Morning Session

9:00 Angelo Vistoli  Twisted curves and logarithmic structures

–9:50

10:00 Constantin Teleman Tensor product K-theory and
–10:50  the stack of G-bundles over a curve
11:00 Coffee
11:30 Pierre Schapira  Quantization-deformation modules on
–12:20  complex symplectic manifolds

Afternoon Session

3:30 Carlos Simpson  TBA

–4:20

4:30 Coffee
5:00 Bertrand Toen  Moduli stack of dg-categories

–5:50

6:00 Jonathan Block  Noncommutative abelian varieties and
–6:50  nonabelian varieties

Friday, July 2, Morning Session

9:00 Anton Alekseev  Deformation quantization of coadjoint orbits

–9:50

10:00 Alberto Cattaneo  Coisotropic submanifolds and dual pairs

–10:50

11:00 Coffee
11:30 Charles-Michel Marle The cotangent groupoid to a semi-direct
–12:20  product of Lie groupoids

Afternoon Session

2:30 Jean-Louis Tu  Twisted K-theory of groupoids

–3:20

3:30 Jouko Mickelsson  Twised K-theory invariants

–4:20

4:30 Coffee